OpenStudy (anonymous):

A = batting average B = times at bat multiplied by player’s seasonal average K = hits in the new game G = times at bat plus times at bat in a new game Solve A = (B+K)/G for K. Please show your work. Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the nearest whole number. Batter A B G K George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (anonymous):

@sami-21 Can You Help Me?

OpenStudy (anonymous):

@experimentX C

OpenStudy (anonymous):

@experimentX Can You Help Me?

OpenStudy (experimentx):

A = (B+K)/G K = AG - B

OpenStudy (anonymous):

Ok..

OpenStudy (experimentx):

George 0.319 10.7 46 K = 0.319*46 - 10.7 = 3.974 do same for every place in K

OpenStudy (anonymous):

Can You Help Me With On One.